# On computing and the complexity of computing higher-order $U$-statistics, exactly

### v1

@@ 3.3 The Algorithm and Time Complexity of Computing $U$-Statistics
Definition 12 (Restricted $V$-statistics).
@@
Lemma 1 (Decomposition of $U$-statistics into $V$-statistics).
@@
where
$\mu_{\pi}=(-1)^{(m-|\pi|)}\prod_{C\in\pi}(|C|-1)!.$
The proof of the above lemma can be found in Appendix B and leverages Möbius inversion from combinatorics. We now discuss how to extend the tensorization algorithm from $V$-statistics (Algorithm 1) to $U$-statistics. Let $\mathbb{U}(h,\bm{X})$ be an $m$-th order $U$-statistic with kernel function $h$ and sample $\bm{X}\in{\mathbb{X}}^{n}$, where $h$ permits a MD structure ${\mathcal{H}}\times{\mathcal{A}}$. On the surface, it seems that we have to apply Algorithm 2 to each resulting $V$-statistics in the decomposition given in Lemma 1. But in fact, we only need to apply Algorithm 2 once since all the resulting $V$-statistics can be tensorized by the same tensor tuple ${\mathcal{T}}=\textsc{Tensorizaion}({\mathcal{H}},\bm{X})$.
@@
$\mathbb{V}[\pi](h,\bm{X})=\mathsf{Einsum}({\mathcal{T}},{\mathcal{A}}_{\pi}).$ (2)
@@
2:An Einsum notation ${\mathcal{A}}=(A_{k})_{k=1}^{K}$ with index set $[m]$
3:A partition $\pi\in\Pi_{m}$ of size $J$
4:The Einsum notation ${\mathcal{A}}_{\pi}$ induced by $\pi$ on $\mathcal{A}$}
5:function InducedNotation(${\mathcal{A}};\pi$)
6:  Let $(Q_{j})_{j=1}^{J}$ be a permutation of $\pi$ $\triangleright$ Label subsets by $j\in[J]$
7:  Define $\psi\colon[m]\to[J]$ as the subset assignment map:
8:  function $\psi$(ii)
9:   return $j$ if $i\in Q_{j}$ $\triangleright$ Map all indices in $Q_{j}$ to $j$
10:  end function
11:  Initialize ${\mathcal{A}}_{\pi}\leftarrow{\mathcal{A}}$
12:  for $k\in[K]$ do
13:   for $i\in[|A_{k}|]$ do
14:     ${\mathcal{A}}_{\pi}[k][i]\leftarrow\psi(A_{k}[i])$ $\triangleright$ Replace index with its subset label
@@
16:  end for
17:  return ${\mathcal{A}}_{\pi}$
18:end function
@@
Then it holds that
$\mathbb{U}({\mathcal{T}},{\mathcal{A}})=\mathbb{U}(\widetilde{{\mathcal{T}}},{\mathcal{A}}).$
@@
With the sparsification trick, the final algorithm of computing $U$-statistics is displayed in Algorithm 4 below.
@@
6:  Let $\widetilde{{\mathcal{T}}}$ be constructed as in Lemma 2
7:  Initialize $u\leftarrow 0$
8:  for $\pi\in\Pi_{m}$ do
9:   if $\forall Q\in\pi,\forall A\in{\mathcal{A}};|Q\cap\mathsf{set}(A)|<2$ then
10:     $\mu_{\pi}\leftarrow(-1)^{m-|\pi|}\prod_{C\in\pi}(|C|-1)!$ $\triangleright$ See Lemma 1
11:     ${\mathcal{A}}_{\pi}\leftarrow$ InducedNotation(${\mathcal{A}};\pi$) $\triangleright$ See Algorithm 3
12:     $v_{\pi}\leftarrow\mathsf{Einsum}(\widetilde{{\mathcal{T}}},{\mathcal{A}}_{\pi})$ (see footnote 1)
13:     $u\leftarrow u+\mu_{\pi}v_{\pi}$
14:   end if
15:  end for
16:  return $u$
17:end function
@@
$-\mathbb{V}[\{1,4\}\{2\}\{3\}](\widetilde{{\mathcal{T}}},{\mathcal{A}})-\cancel{\mathbb{V}[\{2,3\}\{1\}\{4\}](\widetilde{{\mathcal{T}}},{\mathcal{A}})}-\mathbb{V}[\{2,4\}\{1\}\{3\}](\widetilde{{\mathcal{T}}},{\mathcal{A}})-\cancel{\mathbb{V}[\{3,4\}\{1\}\{2\}](\widetilde{{\mathcal{T}}},{\mathcal{A}})}$
$+\cancel{\mathbb{V}[\{1,2\}\{3,4\}](\widetilde{{\mathcal{T}}},{\mathcal{A}})}+\mathbb{V}[\{1,3\}\{2,4\}](\widetilde{{\mathcal{T}}},{\mathcal{A}})+\cancel{\mathbb{V}[\{1,4\}\{2,3\}](\widetilde{{\mathcal{T}}},{\mathcal{A}})}$
$+\cancel{2\mathbb{V}[\{1,2,3\}\{4\}](\widetilde{{\mathcal{T}}},{\mathcal{A}})}+\cancel{2\mathbb{V}[\{1,2,4\}\{3\}](\widetilde{{\mathcal{T}}},{\mathcal{A}})}+\cancel{2\mathbb{V}[\{1,3,4\}\{2\}](\widetilde{{\mathcal{T}}},{\mathcal{A}})}$
$+\cancel{2\mathbb{V}[\{2,3,4\}\{1\}](\widetilde{{\mathcal{T}}},{\mathcal{A}})}-\cancel{\mathbb{V}[\{1,2,3,4\}](\widetilde{{\mathcal{T}}},{\mathcal{A}})}.$
As shown in the above display, there were originally 15 $V$-statistics in total, whereas only 5 $V$-statistics are left after applying the sparsification trick! Taking the second term $\mathbb{V}[\{1,2\}\{3\}\{4\}](\widetilde{{\mathcal{T}}},{\mathcal{A}})$ on the RHS of the above display as an example: $\mathbb{V}[\{1,2\}\{3\}\{4\}](\widetilde{{\mathcal{T}}},{\mathcal{A}})\equiv 0$ because the intersection between $Q=\{1,2\}$ and $A_{1}=(1,2)$ is 2, violating the criterion described in (4).
Armed with the above preparations, we can bound the time complexity of exactly computing $U$-statistics using Proposition 1. Given a partition $\pi$ and the decomposition signature $\mathcal{A}$}, the decomposition graph corresponding to induced Einsum notation ${\mathcal{A}}_{\pi}$ can be characterized by the quotient graph $G_{\mathcal{A}}/\pi$ (see Definition 8) of the decomposition graph $G_{\mathcal{A}}$ induced by $\pi$.
@@
$M$ $\coloneqq\max\{\mathsf{tw}(G_{\mathcal{A}}/\pi)\mid\pi\in\Pi_{m}^{\mathcal{A}}\},$
$N_{l}$ $\coloneqq|\{\pi\in\Pi_{m}^{{\mathcal{A}}}\mid\mathsf{tw}(G_{{\mathcal{A}}}/\pi)=l\}|,\;\forall\,l\in[M].$
@@
In the above result, quantities $N_{l}$ and $M$ are introduced. $N_{l}$ is the number of partitions in $\Pi_{m}^{{\mathcal{A}}}$ such that the treewidth of the decomposition graph $G_{{\mathcal{A}}}$ quotient by $\pi$ is $l$, is defined. Although abstract, $N_{l}$ and $M$ can in principle be computed for concrete examples. For instance, in Section 4.1, we report the values of $M$ for HOIF estimators between orders $2$ to $12$; see the last column of Table 1. At this moment, we are not yet aware of a good strategy to count $N_{l}$, except a loose upper bound by the total number of $V$-statistics in the decomposition; see the third column of Table 1.
@@
In this paper, we have used the treewidth of the decomposition graph (or its quotient as in Corollary 2) to characterize the optimistic time-complexity of exactly computing UU/$V$-statistics, and the Einsum operation. Although treewidth has been a widely used concept for establishing the time-complexity of algorithms on graphs (dong2021nearly), it is in general nontrivial to compute or derive a tight bound. To our knowledge, Theorem 4.7 in (kneis2009bound) proved a bound for treewdith linear in the edge size. We do not know if this bound can be further improved. We will discuss this as an open problem at the end of this paper (Section 5).
Before moving on to applications of our algorithm in the next section, we summarize the steps of exactly computing a $U$-statistic using our proposed algorithm. Given a $U$-statistic kernel $h$ and a sample $\bm{X}$:
@@ 4 Examples and Numerical Results
In this section, we evaluate the performance of our new algorithm for computing UU/$V$-statistics exactly in three concrete statistical applications, including computing higher-order influence functions (Section 4.1), computing motif counts in random graphs (Section 4.2) and computing distance covariance measures (Section 4.3).
As mentioned, we have made our algorithm available for public use as a Python package u-stats, which is available to download from either its GitHub repository (clicking u-stats) or the Python package index PyPi by typing the following on the command line:
pip install u-stats
We also refer interested readers to the GitHub link for how to implement u-stats in practice.
@@ 4.1 Application 1: Higher-Order Influence Functions
As mentioned in the Introduction, Higher-Order Influence Functions (HOIFs) (robins2008higher; robins2016technical) can help construct rate-optimal estimators for important parameters such as the treatment-specific mean under ignorability under certain complexity-reducing assumptions on the data generating mechanism. However, estimators constructed via HOIFs are high-order $U$-statistics, which are time-consuming to compute in practice. Fortunately, the HOIF estimator of ATE is a sum of $U$-statistics, each of which has kernel satisfying the MD structure as in Definition 9. In the setting of an unconfounded observational study, we often have access to a sample tuple $\bm{X}\equiv(X_{i}\equiv(Z_{i},A_{i},Y_{i}))_{i=1}^{n}$ of size $n$, where $Z\in{\mathbb{R}}^{d}$ denotes the baseline covariates, $A\in\{0,1\}$ denotes a binary treatment variable and $Y$ denotes the outcome of interest. According to robins2008higher, with a $k$-dimensional dictionary $\phi:{\mathbb{R}}^{d}\rightarrow{\mathbb{R}}^{k}$, an $m$-th order influence function estimator of the treatment-specific mean reads as
@@
$h^{(j)}(\bm{X}[\bar{i}_{j}])=A_{i_{1}}\phi(Z_{i_{1}})^{\top}\left\{\prod_{s=3}^{j}\left(A_{i_{s}}\phi(Z_{i_{s}})A_{i_{s}}\phi(Z_{i_{s}})^{\top}-\mathbb{I}\right)\right\}\phi(Z_{i_{2}})Y_{i_{2}}$
@@
Then obviously $h_{j}^{(m)}$ permits a MD structure by letting $f_{l}(X_{1},X_{2})\coloneqq A_{1}\phi(Z_{1})^{\top}A_{2}\phi(Z_{2})$ for $l<j$ and $f_{j}(X_{1},X_{2})\coloneqq A_{1}\phi(Z_{1})^{\top}\phi(Z_{2})Y_{2}$. The corresponding decomposition signature of $h_{j}^{(m)}$ is $((i,i+1))_{i=1}^{j-1}$, which is denoted by ${\mathcal{A}}_{\mathsf{HOIF},j}$, so the decomposition graph of ${\mathcal{A}}_{\mathsf{HOIF},j}$ is a chain graph with e=j1e=j-1 edges.
We also demonstrate the effectiveness of the sparsification trick using the HOIF example (see Lemma 2). The following table enlists the total number of $V$-statistics and the number of floating-point operations (FLOPs) that need to be computed before and after applying the sparsification trick when the sample size is $10^{4}$. The rightmost column of Table 1 also records the value of $M$ defined in Corollary 2. The derivation of this column can be found in Appendix C.2 (and unfortunately we do not know a more systematic way of deriving these).
Table 1: Exact number of floating-point operations (FLOPs), $V$-statistic decompositions, and intermediate tensor sizes across HOIF orders from $2$ to $12$, with and without the sparsification trick, where the Einsum ordering is found by applying opt-einsum with optimization scheme greedy. Sample size is fixed at $n=10000$.
Order ($m$)
V-stat terms
(Bell number)
V-stat terms
FLOPs
FLOPs
$M$
2 2 1 $2.00020\times 10^{8}$ $2.00000\times 10^{8}$ 11
3 5 2 $060\times 10^{12} 2.00020$$020\times 10^{12}$ 11
4 15 5 $8.00310\times 10^{12}$ $4.00150\times 10^{12}$ 22
5 52 15 $5.20133\times 10^{13}$ $60045\times 10^{13} 22$
6 203 52 $2.94069\times 10^{14}$ $1.20020\times 10^{14}$ 22
7 877 203 $1.78236\times 10^{15}$ $6.32091\times 10^{14}$ 22
8 4140 877 $31054\times 10^{17} 2.2$$23430\times 10^{17}$ 33
9 21147 4140 $5.29126\times 10^{18}$ $47959\times 10^{18} 33$
10 115975 21147 $2.07990\times 10^{21}$ $2.04048\times 10^{21}$ 33
11 678570 115975 $5.50026\times 10^{22}$ $3.72144\times 10^{22}$ 44
12 4213597 678570 $1.13401\times 10^{24}$ $6.62181\times 10^{23}$ 44
In Tables 2 and 3, we report the average runtime over 3 runs when computing the $U$-statistic $\mathbb{U}(h_{j}^{(m)},\bm{X})$ where kernel function $h_{j}^{(m)}$ is as in (5) with $m=7$ and $j\in\{2,3,\cdots,7\}$ and sample size $n\in\{1000,2000,4000,8000,10000\}$, respectively using CPU parallel computation and using a single GPU. We would like to highlight that even when $m=7$, $n=10000$, it only takes about 5 minutes using our new package u-stats, with CPU parallel computation. If one instead has access to a single GPU, the runtime even reduces to about 12 seconds, a substantial speedup! We remark in passing that all the runtime reported in Table 2 and Table 3 excludes the time of kernel evaluation and tensorization.
Table 2: Average runtime (in seconds) using CPU parallel computation. Experiments were conducted on Intel Xeon ICX Platinum 8358 CPUs (2.6GHz, 64 total cores) with memory of 512 GB, evaluated across varying sample sizes and orders of HOIF-type $U$-statistics.
@@
2 0.64 0.00141 0.01298 0.03174 0.04746
3 0.00396 0.01518 0.07392 0.26849 0.45976
4 0.02321 0.07313 0.32765 2.09545 2.36766
5 0.09853 0.36239 1.71419 9.11349 14.21350
6 0.38878 1.47719 7.44444 40.22143 58.85063
7 1.91677 6.75947 34.08805 195.31414 290.54295
Table 3: Average runtime (in seconds) using a single GPU (NVIDIA RTX 4090, 24GB) with parallel computation, across varying sample sizes and orders of HOIF-type $U$-statistics.
Order 1000 2000 4000 8000 10000
2 0.00184 0.00151 0.00155 0.00238 0.00339
3 0.00172 0.00147 0.00193 0.00577 0.00745
4 0.00305 0.00353 0.00707 0.03612 0.06245
5 0.00835 0.01089 0.03456 0.19807 0.36160
6 0.02608 0.03906 0.16981 1.12072 2.13328
7 0.12031 0.19132 0.91884 6.45225 12.21350
4.2 Application 2: Motif Counts in Random Graph Models and Networks
Motif (or subgraph) counts refer to the number of occurrences of a particular subgraph type $\mathsf{r}$ within a larger graph $G$, denoted by $\mathsf{C}(\mathsf{r},G)$. These counts, also known as network moments, are crucial for understanding the structural properties of random graph models often used to model real-world networks. Interestingly, motif counts can be equivalently represented as $U$-statistics. Leveraging this observation, we show that various motif counts can be computed using a single Einsum operation. While this approach may not capture fine-grained structural information such as sparsity of the adjacency matrix, it is highly efficient in the context of dense graphs, significantly outperforming a popular software called igraph designed for computing motif counts.

### v2

@@ 3.3 The Algorithm of Computing $U$-Statistics
Definition 12 (Restricted $V$-Statistic).
@@
Lemma 1 (Decomposition of $U$-Statistic into $V$-Statistics).
@@
where $\mu_{\pi}=(-1)^{(m-|\pi|)}\prod_{C\in\pi}(|C|-1)!$.
[…]
@@
$\mathbb{V}[\pi](h,\bm{X})=\textsc{Einsum}(\mathcal{T},\mathcal{A}_{\pi}).$ (2)
@@
2:An Einsum notation $\mathcal{A}=(A_{k})_{k=1}^{K}$ with index set $[m]$ and a partition $\pi\in\Pi_{m}$ of size $J$
3:The Einsum notation $\mathcal{A}_{\pi}$ induced by $\pi$ on $\mathcal{A}$
4:function InducedNotation($\mathcal{A};\pi$)
5:  Let $(Q_{j})_{j=1}^{J}$ be a permutation of $\pi$ $\triangleright$ Label subsets by $j\in[J]$
6:  Define $\psi\colon[m]\to[J]$ as the subset assignment map:
7:  function $\psi$(ii)
8:   return $j$ if $i\in Q_{j}$ $\triangleright$ Map all indices in $Q_{j}$ to $j$
9:  end function
10:  Initialize $\mathcal{A}_{\pi}\leftarrow\mathcal{A}$
11:  for $k\in[K]$ do
12:   for $i\in[|A_{k}|]$ do
13:     $\mathcal{A}_{\pi}[k][i]\leftarrow\psi(A_{k}[i])$ $\triangleright$ Replace index with its subset label
14:   end for
@@
16:  return $\mathcal{A}_{\pi}$
17:end function
@@
Then $\mathbb{U}(\mathcal{T},\mathcal{A})=\mathbb{U}(\widetilde{\mathcal{T}},\mathcal{A})$.
@@
Finally, we present the new algorithm for computing $U$-statistics in Algorithm 4 below.
@@
6:  Let $\widetilde{\mathcal{T}}$ be constructed as in Lemma 2 and initialize $u\leftarrow 0$
7:  for $\pi\in\Pi_{m}$ do
8:   if $\forall Q\in\pi,\forall A\in\mathcal{A};|Q\cap\mathsf{set}(A)|<2$ then
9:     $\mu_{\pi}\leftarrow(-1)^{m-|\pi|}\prod_{C\in\pi}(|C|-1)!$ $\triangleright$ See Lemma 1
10:     $\mathcal{A}_{\pi}\leftarrow$ InducedNotation($\mathcal{A};\pi$) $\triangleright$ See Algorithm 3
11:     $v_{\pi}\leftarrow\textsc{Einsum}(\widetilde{\mathcal{T}},\mathcal{A}_{\pi})$ $\triangleright$ See Footnote 1
12:     $u\leftarrow u+\mu_{\pi}v_{\pi}$
13:   end if
14:  end for
15:  return $u$
16:end function
@@
$-\mathbb{V}[\{1,4\}\{2\}\{3\}](\widetilde{\mathcal{T}},\mathcal{A})-\cancel{\mathbb{V}[\{2,3\}\{1\}\{4\}](\widetilde{\mathcal{T}},\mathcal{A})}-\mathbb{V}[\{2,4\}\{1\}\{3\}](\widetilde{\mathcal{T}},\mathcal{A})$
$-\cancel{\mathbb{V}[\{3,4\}\{1\}\{2\}](\widetilde{\mathcal{T}},\mathcal{A})}+\cancel{\mathbb{V}[\{1,2\}\{3,4\}](\widetilde{\mathcal{T}},\mathcal{A})}+\mathbb{V}[\{1,3\}\{2,4\}](\widetilde{\mathcal{T}},\mathcal{A})$
$+\cancel{\mathbb{V}[\{1,4\}\{2,3\}](\widetilde{\mathcal{T}},\mathcal{A})}+\cancel{2\mathbb{V}[\{1,2,3\}\{4\}](\widetilde{\mathcal{T}},\mathcal{A})}+\cancel{2\mathbb{V}[\{1,2,4\}\{3\}](\widetilde{\mathcal{T}},\mathcal{A})}$
$+\cancel{2\mathbb{V}[\{1,3,4\}\{2\}](\widetilde{\mathcal{T}},\mathcal{A})}+\cancel{2\mathbb{V}[\{2,3,4\}\{1\}](\widetilde{\mathcal{T}},\mathcal{A})}-\cancel{\mathbb{V}[\{1,2,3,4\}](\widetilde{\mathcal{T}},\mathcal{A})}.$
As shown in the above display, there were originally 15 non-redundant $V$-statistics in total, whereas only 5 $V$-statistics remain after applying the sparsification trick!
Taking the second term $\mathbb{V}[\{1,2\}\{3\}\{4\}](\widetilde{\mathcal{T}},\mathcal{A})$ on the RHS of the above display as an example: $\mathbb{V}[\{1,2\}\{3\}\{4\}](\widetilde{\mathcal{T}},\mathcal{A})\equiv 0$ because the intersection between $Q=\{1,2\}$ and $A_{1}=(1,2)$ is 2, violating the criterion described in (4).
Given a partition $\pi$ and the decomposition signature $\mathcal{A}$, the decomposition graph corresponding to induced Einsum notation $\mathcal{A}_{\pi}$ can be characterized by the quotient graph $G_{\mathcal{A}}/\pi$ (see Definition 8) of the decomposition graph $G_{\mathcal{A}}$ induced by $\pi$. Armed with the above observation, we can bound the time complexity of exactly computing $U$-statistics by combining Corollary 1, Lemma 1, and Lemma 2, leading to the following result.
@@
$M\coloneqq\max\{\mathsf{tw}(G_{\mathcal{A}}/\pi)\mid\pi\in\Pi_{m}^{\mathcal{A}}\},\quad N_{l}\coloneqq|\{\pi\in\Pi_{m}^{\mathcal{A}}\mid\mathsf{tw}(G_{\mathcal{A}}/\pi)=l\}|,\;\forall\,l\in[M].$
@@
[…]
@@
[…]
Before moving on to applications of our algorithm in the next section, we summarize our proposed algorithm below. Given a $U$-statistic kernel $h$ and a sample $\bm{X}$:
@@ 4 Examples and Numerical Results
In this section, we evaluate the performance of our new algorithm for computing UU/$V$-statistics in three concrete statistical applications [^2] The replication code can be found in this link.. We have made our algorithm available for public use, which is available to download from the GitHub repository (by clicking u-stats or its R interface) or the Python package index PyPi. For the first application (Section 4.1) on Higher-Order Influence Functions (HOIFs) (robins2008higher), we also provide a standalone implementation based on u-stats, given the growing interest in using HOIFs in practice in the causal inference community (wager2024causal; zheng2025perturbed; cinelli2025challenges).
@@ 4.1 Application 1: Higher-Order Influence Functions
As mentioned in the Introduction, HOIFs (robins2008higher) can help construct rate-optimal estimators for important parameters such as the treatment-specific mean under ignorability under certain complexity-reducing assumptions on the data generating mechanism. However, estimators constructed via HOIFs are high-order $U$-statistics, which are time-consuming to compute in practice. Fortunately, the HOIF estimator of ATE is a sum of $U$-statistics, each of which has kernel satisfying the MD structure as in Definition 9. In the setting of an unconfounded observational study, we often have access to a sample tuple $\bm{X}\equiv(X_{i}\equiv(Z_{i},A_{i},Y_{i}))_{i=1}^{n}$ of size $n$, where $Z\in{\mathbb{R}}^{d}$ denotes the baseline covariates, $A\in\{0,1\}$ denotes a binary treatment variable and $Y$ denotes the outcome of interest. According to robins2008higher, with a $k$-dimensional dictionary $\phi:{\mathbb{R}}^{d}\rightarrow{\mathbb{R}}^{k}$, an $m$-th order influence function estimator of the treatment-specific mean reads as
@@
$h^{(j)}(\bm{X}[\bar{i}_{j}])=A_{i_{1}}\phi(Z_{i_{1}})^{\top}\left\{\prod_{s=3}^{j}\left(A_{i_{s}}\phi(Z_{i_{s}})\phi(Z_{i_{s}})^{\top}-\mathbb{I}\right)\right\}\phi(Z_{i_{2}})Y_{i_{2}}$
@@
[…]
In Table 1, we report the average runtime over 10 runs when computing the $U$-statistic $\mathbb{U}(h_{j}^{(m)},\bm{X})$ where kernel function $h_{j}^{(m)}$ is as in (5) with $m=7$ and $j\in\{2,3,\cdots,7\}$ and sample size $n\in\{1000,2000,4000,8000,10000\}$, respectively using CPU parallel computation and using a single GPU. We would like to highlight that even when $m=7$, $n=10000$, it only takes about 5 minutes using our new package u-stats, with CPU parallel computation. If one instead has access to a single GPU, the runtime even reduces to about 8 seconds, a substantial speedup! We remark in passing that all the runtime reported in Table 1 excludes the time of kernel evaluation and tensorization.
Table 1: Average runtime (in seconds) using CPU (Intel Xeon ICX Platinum 8358 CPUs [2.6GHz, 64 total cores] with memory of 512 GB) with parallel computation (upper cell) and using a single GPU (NVIDIA RTX 5090, 32GB) with parallel computation (lower cell), evaluated across varying sample sizes and orders of HOIFs.
@@
2 0.00217 0.00070 0.00515 0.00142 0.01965 0.00492 0.03466 0.02779 0.04811 0.04398
3 0.00388 0.00124 0.01548 0.00274 0.06968 0.01337 0.27130 0.06413 0.43110 0.08977
4 0.01963 0.00242 0.07295 0.00475 0.32076 0.02375 1.81642 0.10690 2.33967 0.17264
5 0.09356 0.00567 0.36376 0.00922 1.91751 0.04619 9.80295 0.23752 16.92521 0.41339
6 0.45275 0.01863 1.52683 0.02545 7.76946 0.13763 41.78568 0.89001 62.54028 1.66450
7 1.73287 0.07340 6.80758 0.10322 36.41280 0.64650 197.05856 4.49332 288.89672 8.58869
4.2 Application 2: Motif Counts in Networks
Motif (or subgraph) counts refer to the number of occurrences of a particular subgraph type $\mathsf{r}$ within a larger graph $G$, denoted by $\mathsf{C}(\mathsf{r},G)$. These counts, also known as network moments, are crucial for understanding the structural properties of random graph models often used to model real-world networks. Interestingly, motif counts can be equivalently represented as $U$-statistics. Leveraging this observation, we show that various motif counts can be computed using a single Einsum operation. While this approach may not capture fine-grained structural information such as sparsity of the adjacency matrix, it is highly efficient in the context of dense graphs.